OpenStudy (saadi101):

Solve using substitution: x - 4y = 4 and 2x - 3y = -3. Put answer in coordinate form.

OpenStudy (danjs):

substitution is reducing thenumber of variables in an equation.. one equation with 2 variables does not have a unique solution you can figure one variable in terms of the other..

OpenStudy (danjs):

solve one of the two equations for x or for y the y value is the same in the solution, so you can use what y= and replace that in the other equation

OpenStudy (danjs):

ill solve the first one for x x - 4y = 4, so x = 4y + 4 with that, you can replace x in the other equation by it's value here 4y+4 2x - 3y = -3 ---> 2*(4y + 4) - 3y = -3 there you have a single variable in the equation, and can solve for a value of y,

OpenStudy (mathmale):

Dan, for the third time, PLEASE involve your student in finding his or her answer, and do NOT just hand out yours.